# LLM Fingerprinting via Semantically Conditioned Watermarks

### v2

@@ E.1 Watermark Tokens as Domain Trigger
In Fig.˜8 (left), we plot the ROC curve of this experiment, and the identity line in gray as a reference. We see that the model almost perfectly learns the trigger, outputting a strong watermark when $t_{w}$ is present (above 95% TPR at 1%), while outputting almost no watermarked text in the absence of $t_{w}$. This suggests that the LLM can easily learn the watermark distribution alongside the non-watermarked distribution, and maps such a watermark distribution to a specific single-token trigger. While in Section˜5 we only consider domains as triggers, one could use any token combination as a trigger, improving the harmlessness of the fingerprint at the cost of stealthiness, while benefiting from the strong reliability and guarantees of watermarking-based fingerprints.
@@
In Fig.˜8 (right), we plot the ROC curves across all datasets, as well as the identity line for reference. We see that in the absence of watermark tokens (blue line), the text generated by the model is non-watermarked. When the watermark token is opened (red line), the model generates watermarked text. Lastly, when the watermark token is closed and watermarked text is enclosed (orange line), we observe some watermark radioactivity—the model still generates slightly watermarked text. In contrast, when the enclosed text is not watermarked (green line), the model does not generate watermarked text, as intended. These results show how flexible semantically conditioned watermarking can be. Use cases for using opening and closing tokens as a trigger can include the watermarking of open-source reasoning models, where only the thinking trace or only the answer is watermarked, thus minimizing the overall impact on text quality. We leave this direction for future work.
@@ E.2 Watermarking Multiple Token-Conditioned Domains
In Fig.˜9, we show the ROC curves, running the detection with every key for each scenario. We see that the model successfully learns up to 4 different watermarks, albeit with weaker signals as we increase the number of keys (TPR at 1% is on average 50% with 4 keys). Moreover, we see that even with the detector corresponding to another key (for instance, detector 1 with [WM2] instead of [WM0]), the detection rate is sometimes abnormally high. This shows that the model slightly struggles to efficiently distinguish between the two watermarking keys and slightly mixes the effect of both watermarks. Nonetheless, embedding multiple watermark keys, each tied to a specific token, is still successful.
@@ E.3 Watermarking Other Domains (Harmful Content)
In Fig.˜10, we show the ROC curves for the harmful domain and a general domain (Alpaca). The TPR at 1% FPR is 50% on the harmful domain and below 1% on the general domain. This means that even when jailbroken, the model still outputs watermarked text. These results align with previous work on self-identification of harmful data (Xhonneux et al., 2025), where a model is trained to add a detectable signal when and only when generating harmful content.

### v3

@@ E.1 Watermark Tokens as Domain Trigger
In Fig.˜10 (left), we plot the ROC curve of this experiment, and the identity line in gray as a reference. We see that the model almost perfectly learns the trigger, outputting a strong watermark when $t_{w}$ is present (above 95% TPR at 1%), while outputting almost no watermarked text in the absence of $t_{w}$. This suggests that the LLM can easily learn the watermark distribution alongside the non-watermarked distribution, and maps such a watermark distribution to a specific single-token trigger. While in Section˜5 we only consider domains as triggers, one could use any token combination as a trigger, improving the harmlessness of the fingerprint at the cost of stealthiness, while benefiting from the strong reliability and guarantees of watermarking-based fingerprints.
@@
In Fig.˜10 (right), we plot the ROC curves across all datasets, as well as the identity line for reference. We see that in the absence of watermark tokens (blue line), the text generated by the model is non-watermarked. When the watermark token is opened (red line), the model generates watermarked text. Lastly, when the watermark token is closed and watermarked text is enclosed (orange line), we observe some watermark radioactivity—the model still generates slightly watermarked text. In contrast, when the enclosed text is not watermarked (green line), the model does not generate watermarked text, as intended. These results show how flexible semantically conditioned watermarking can be. Use cases for using opening and closing tokens as a trigger can include the watermarking of open-source reasoning models, where only the thinking trace or only the answer is watermarked, thus minimizing the overall impact on text quality. We leave this direction for future work.
@@ E.2 Watermarking Multiple Token-Conditioned Domains
In Fig.˜11, we show the ROC curves, running the detection with every key for each scenario. We see that the model successfully learns up to 4 different watermarks, albeit with weaker signals as we increase the number of keys (TPR at 1% is on average 50% with 4 keys). Moreover, we see that even with the detector corresponding to another key (for instance, detector 1 with [WM2] instead of [WM0]), the detection rate is sometimes abnormally high. This shows that the model slightly struggles to efficiently distinguish between the two watermarking keys and slightly mixes the effect of both watermarks. Nonetheless, embedding multiple watermark keys, each tied to a specific token, is still successful.
@@ E.3 Watermarking Other Domains (Harmful Content)
In Fig.˜12, we show the ROC curves for the harmful domain and a general domain (Alpaca). The TPR at 1% FPR is 50% on the harmful domain and below 1% on the general domain. This means that even when jailbroken, the model still outputs watermarked text. These results align with previous work on self-identification of harmful data (Xhonneux et al., 2025), where a model is trained to add a detectable signal when and only when generating harmful content.
@@ F.2 Resources
Table 9: Cost Comparison We compare the number of tokens needed to embed the fingerprint into the model for our approach and the two baselines (IF and scalable). The reported prices are based on Lambda Labs pricing.
IF Scalable Ours
# tokens $\sim$0.71M $\sim$7.19M $\sim$81.92M
Cost \$<1.0<1.0 \$<1.0<1.0 $\sim$\$$6.0$
@@
Table˜9 shows the number of tokens needed for embedding the different fingerprints. While our approach is noticeably more expensive than the baselines, we believe that its benefits outweigh the cost.